OpenStudy (anonymous):

WILL HAPPILY GIVE A MEDAL. I LOVE YOU GUYS. Jesse's condo has a length 10 feet longer than the width. If his condo is 1,064 square feet, what are the dimensions? A. 28 feet by 38 feet B. 19 feet by 56 feet C. 14 feet by 76 feet D. 32.6 feet by 32.6 feet

OpenStudy (anonymous):

OKay, so the first thing you need to do is set up the equations the word problem gives you :) So, you have L = W + 10 -> Because the Length is 10 feet longer the WIdth and then the equation for area, which is L * W = Area

OpenStudy (anonymous):

With that, you substitute the equation for L into the equation for Area. So, Area = (W+10) * W And then plug in the area that you know 1064 = (W+10) * W and Solve for W :) Do you know how to do that?

OpenStudy (anonymous):

It's A!

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Great job!! You're very welcome! :D